# Bridging the Gap Between Saliency Prediction and Image Quality Assessment thanks: This study was supported by Russian Science Foundation under grant 24-21-00172, https://rscf.ru/en/project/24-21-00172/

### v1

@@ Abstract
Over the past few years, deep neural models have made considerable advances in image quality assessment (IQA). However, the underlying reasons for their success remain unclear, owing to the complex nature of deep neural networks. IQA aims to describe how the human visual system (HVS) works and to create its efficient approximations. On the other hand, Saliency Prediction task aims to emulate HVS via determining areas of visual interest. Thus, we believe that saliency plays a crucial role in human perception.
In this work, we conduct an empirical study that reveals the relation between IQA and Saliency Prediction tasks, demonstrating that the former incorporates knowledge of the latter. Moreover, we introduce a novel SACID dataset of saliency-aware compressed images and conduct a large-scale comparison of classic and neural-based IQA methods. Supplementary code and data will be available at https://github.com/Alexkkir/SACID.
@@ I Introduction
Image Quality Assessment (IQA) aims to estimate perceptual image quality similar to the human visual system (HVS). It can be useful in enhancing the visual experience of humans in various applications such as image compression, restoration, acquisition, editing, and generation.
Despite IQA models achieving performance similar to human assessors on various benchmarks, the inner mechanisms of IQA models remain unclear. A question of interest is whether these IQA models implicitly incorporate characteristics from the human visual system (HVS), especially in focusing attention on certain image regions, a principle known as saliency [1, 2]. Saliency Prediction (SP), as a specialized field, has seen the emergence of SOTA models capable of determining the points of human visual interest.
Figure 1: Saliency and GradCAM identify important regions of image. Upper row – saliency maps, predicted with TranSalNet[3]. Bottom row – GradCAMs, build for Baseline-EfficientB0.
@@
  • We propose a methodology to extract saliency maps from trained IQA models, which reveals that learning-based IQA methods incorporate saliency in their predictions and can outperform strong SP baselines like center-prior.
@@ II-A Image Quality Assessment
IQA aims to measure the perceptual quality of an image by assessing the extent of image artifacts and distortions. According to the requirement for ground-truth reference images, most IQA techniques can be categorized as Full-Reference (FR) or No-Reference (NR). The latter scenario is more relevant for real-world applications [4].
Early approaches based on natural scene statistics (NSS) [5, 6, 7, 8, 9], employed hand-crafted features and a regression model above them to predict the mean opinion score (MOS). However, as deep learning has evolved, neural network-based approaches have become popular. Usually, they consist of a convolutional or transformer backbone, trained end-to-end with a regression head. Many studies have suggested iterative improvements of these architectures. HyperIQA [10] proposed a multiscale feature; MUSIQ [11] employed multiscale inputs and a vision transformer backbone to process images with varying sizes and aspect ratios.
Several authors [12, 13] have proposed methods for creating paired models that can serve in both FR and NR situations. For example, TOPIQ utilizes features from the original, distorted, and difference images at each layer.
@@ II-B Saliency in Image Quality Assessment
Prior research indicates a connection between saliency and Image Quality Assessment [12, 14, 15, 16, 17, 18, 19, 20, 21 ]. Efforts are underway to enhance IQA-model performance by incorporating saliency [15, 22, 23, 24, 25, 26, 27, 28]. Early studies [9] have demonstrated improvement in NSS-based metrics like PSNR and SSIM by reweighting the error map based on saliency. In SGDNet [22], authors introduced a separate head to predict saliency and then applied its predictions to reweight features from the final layer.
TranSLA [24] incorporates saliency as a query branch in cross-attention, while HVS-5M [23] employs a pre-trained SP model to reweight features from the final layer. In SCVC [26], authors split the image into patches and aggregated their scores using a Gaussian function and saliency maps (SM). Furthermore, according to [12], LPIPS [29] performance can be enhanced by spatially reweighting feature maps. However, several studies [14] report that modern video quality assessment models (VQA) face only slight improvement from using saliency. Notably, our research reveals that models already possess knowledge of saliency, enabling them to incorporate it into their predictions.
@@ II-C AI Explainability
Even though neural networks reached a human level performance in numerous tasks, they remain a black box. Special methods thus have been developed to provide insight into their internal operations. The first effort for explaining CNNs, CAM [30], considers a simple case of a convolutional network with a head consisting of a single linear layer after global average pooling (GAP). To obtain attention maps, features from the last layer are summed with weights from the fully connected layer. GradCAM [31] extends CAM to models with arbitrary head, employing gradients w.r.t. the model’s predictions rather than relying on weights. Subsequent works [32, 33, 34] propose heuristics for improving the method.
In another vein, some researchers have studied benchmarking explanation maps. In [32] they suggest masking the most important image areas and monitoring changes in model confidence. The authors of [35] retrained the model on perturbed images to reduce domain shift. Work [36] introduced a metric, area over the perturbation curve, as a measure of explanation-map fidelity. In [37] the researchers propose a novel method for masking image areas.
Furthermore, several works address explainability in IQA. For instance, in [13] authors showed that IQA models only require half of an image to make accurate predictions. To identify essential areas, they divided each image into 12 square patches and examined predictions of the transformer-based IQA model as they masked their various combinations. Then they measured the accuracy of SOTA IQA model on trivial and important regions that make up half of the area of the original image. They found out that when using the former, the quality decreases, while when using the latter, the quality of the model is almost preserved. We simplify this approach and show that masking pixels based on saliency maps or GradCAM yields comparable results.
@@ III-A Extracting Saliency from IQA models
TABLE I: Performance comparison on SALICON[38] dataset. GradCAM-extracted saliency maps from IQA models outperform center-prior baseline, but are surpassed by SOTA SP methods. The best results are bold, the second best are underlined, and the third best are italics.
@@
We started our experiments by testing the hypothesis that ground-truth saliency (e.g. from eye-tracker) correlates with explanation GradCAM maps of IQA models. In our experiments we used TOPIQ [12], DBCNN [39] and CLIP-IQA [40]. We also trained our simple baseline model, consisting of a backbone, GAP pooling, and MLP head. We tested two backbones: EfficientNet-B0 [42] and ResNet-50 [43]. To build GradCAM maps, we choose the HiResCAM [33] method and applied smoothing through augmentations and SVD decompositions of feature maps as recommended in [44].
We used straightforward saliency baselines – CenterPrior, a Gaussian distribution that approximates the saliency averaged over the dataset – and two SOTA saliency prediction models [41, 3]. We evaluate the model quality based on common saliency metrics – NSS, SIM, CC, and KLD calculated on validation split of SALICON dataset [38]. Before calculating metrics, we employed map transforms [45].
@@ III-B Saliency Masking
Inspired by [13], we employed saliency and GradCAM maps for masking image regions. For better interpretability, we masked different image portions with two strategies: Most Relevant First (MoRF) and Least Relevant First (LeRF) [36]. We applied perturbations such as filling with black color and ImageNet mean. Before thresholding explanation maps, we apply Gaussian blur with a large kernel, to ensure we can calculate all quantiles. Fig. 1 shows examples of images after applying the first perturbation. We evaluated the correlation drop on the KonIQ-10k dataset; Fig. 2 presents the results.
Masking of salient areas decreases quality more than masking of trivial areas does. The same observation holds for GradCAM. Interestingly, masking with saliency achieves a larger performance gap between LeRF and MoRF regions compared to [13], even though the latter employs knowledge of pre-trained IQA models. This finding indicates that saliency addresses the regions that are crucial for IQA models. Furthermore, we note that the behavior of models when masked with GradCAM maps and saliency exhibits significant similarity, which we can interpret as an existing relationship between the attention of IQA models and humans.
@@ III-C Dual-Task Training
After discovering a close connection between IQA and saliency prediction (SP), we decided to train a model on both tasks simultaneously and determine whether we could do so without decreasing performance. Our effort took two approaches.
In our first approach, denoted as Baseline-Sal-Loss, we simply added to our baseline model a small decoder, consisting of a single convolution with a 1×1 kernel followed by a sigmoid function. Our second approach employed the model’s GradCAM as the second output by incorporating it into the loss function. This enabled us to avoid adding any extra parameters to the model. The GradCAM calculation is as follows:
$L^{C}_{\text{GradCAM}}=\operatorname{ReLU}\underbrace{\left(\sum_{k}\alpha^{c}_{k}A^{k}\right)}_{\mathclap{\text{linear imputation}}},\quad\alpha^{c}_{k}=\overbrace{\frac{1}{Z}\sum_{i}\sum_{j}}^{\mathclap{\text{global average pooling}}}\underbrace{\frac{\partial y^{c}}{\partial A^{k}_{ij}}}_{\mathllap{\text{gradients via backprop}}}$
In our case, $c=1$. To ensure differentiability, we detached $\alpha^{c}_{l}$ from the computational graphs and calculated gradients only over channel maps. This resulted in a model called Baseline-GradCAM-Loss.
We used a random 80% of the KonIQ-10k dataset [46] for training and the remaining 20% for testing as well as CLIVE [47], SALICON [38], and CAT2000 [48] datasets. As KonIQ-10k doesn’t provide SM, we used proxy SM generated with MSINet[41]. We repeated each experiment 10 times and presented averaged results in Tab. II. We compared these dual-task models with the baseline model trained solely on IQA (with III-A saliency extraction technique), CenterPrior, and SOTA saliency models. The findings indicate that both approaches allow us to train models on both tasks without decreasing the performance of IQA while significantly improving SP results.
@@ III-D Non-uniform-Compression Dataset
Subsequently, we calculated SROCC, PLCC and KRCC for each metric from the PYIQA toolbox [50], saliency-aware versions of PSNR and SSIM (EW-PSNR and EW-SSIM [9]), our model variants (Baseline, Baseline-Sal-Loss, and Baseline-GradCAM-Loss), and salient deep-learning models (SGDNet[22], etc.). Each image was provided in its original resolution to maximize the performance of most models. KRCC was calculated in groups with different both sequences and bitrates; thus we denote it as $\text{KRCC}^{\dagger}$.
@@
We also observed intriguing disparities regarding to TOPIQ model: TOPIQ-NR achieved notably higher scores than TOPIQ-FR. Moreover, most NR metrics are ranked above FR (as depicted in Fig. 3).
@@ IV Discussion
Type Method SROCC $\uparrow$ PLCC $\uparrow$ $\text{KRCC}^{\dagger}$uparrow
NSS based MS-SSIM [6] 0.807 0.849 0.013
BRISQUE [8] 0.817 0.833 0.144
PSNR 0.822 0.858 0.009
SSIM [5] 0.835 0.880 0.004
VIF [7] 0.848 0.891 0.052
EW-SSIM [9] 0.850 0.867 0.272
EW-PSNR [9] 0.875 0.893 0.356
FR AHIQ [51] 0.789 0.817 0.091
LPIPS [29] 0.818 0.854 0.024
PieAPP [52] 0.833 0.869 0.069
TOPIQ-FR [12] 0.836 0.848 0.257
DISTS [53] 0.871 0.904 0.178
NR MANIQA [54] 0.752 0.773 0.128
TReS [55] 0.765 0.800 0.117
HyperIQA [10] 0.784 0.802 0.124
SGDNet-None [22] 0.820 0.853 0.302
PaQ2PiQ [56] 0.848 0.858 0.237
SGDNet-Output [22] 0.854 0.879 0.374
Baseline-GradCAM-Loss 0.863 0.865 0.248
CLIP-IQA+ [40] 0.869 0.896 0.196
Baseline 0.871 0.881 0.217
TOPIQ-NR [12] 0.876 0.886 0.257
Baseline-Sal-Loss 0.880 0.886 0.259
MUSIQ [11] 0.888 0.899 0.324
DBCNN [39] 0.901 0.917 0.331
Initially, our experiments focused on developing a saliency-aware IQA metric and exploring its potential to enhance IQA performance, given their evident relationship. However, we encountered challenges, prompting a shift in focus to investigate the underlying reasons.
We visualized GradCAMs of several IQA-model and noted their similarity to SM even though the IQA should pay more attention to areas with artifacts. Hence, we hypothesized that learnable models inherently understand saliency patterns. To confirm this, we introduced a method for extracting saliency maps from IQA models, showed that they mostly rely on salient regions, and proved the possibility of training a single model without additional parameters for both tasks without performance degradation.
Figure 3: Correlations of top-10 full-reference and no-referene methods on proposed SACID dataset.
We also questioned the adequacy of existing datasets, as they primarily contain images compressed with conventional codecs. To overcome this limitation, we created a new dataset consisting of nonuniformly compressed images. We tested various IQA metrics on this dataset and confirmed that incorporating saliency can enhance simple metrics such as PSNR and SSIM, however, deep-learning-based failed to see a similar boost in quality. Additionally, we observed that no-reference metrics outperformed full-reference metrics.
As a future work, we are considering improving the robustness of metrics to adversarial attacks by utilizing saliency.

### v2

@@ Abstract
Over the past few years, deep neural models have made considerable advances in image quality assessment (IQA). However, the underlying reasons for their success remain unclear due to the complex nature of deep neural networks. IQA aims to describe how the human visual system (HVS) works and to create its efficient approximations. On the other hand, Saliency Prediction task aims to emulate HVS by determining areas of visual interest. Thus, we believe that saliency plays a crucial role in human perception.
In this work, we conduct an empirical study that reveals the relation between IQA and Saliency Prediction tasks, demonstrating that the former incorporates knowledge of the latter. Moreover, we introduce a novel SACID dataset of saliency-aware compressed images and conduct a large-scale comparison of classic and neural-based IQA methods. Supplementary code and data will be available at https://huggingface.co/datasets/alexkkir/SACID.
@@ I Introduction
Image Quality Assessment (IQA) aims to measure image quality aligned with human visual perception. Improving IQA can greatly enhance user experience in tasks such as image compression, restoration, editing, and generation.
Despite achieving high correlations with human judgments, the internal workings of IQA models remain unclear. An important open question is whether IQA models implicitly adopt properties of human vision, particularly visual saliency—the human tendency to focus attention on certain image regions [1, 2]. Meanwhile, Saliency Prediction (SP) has advanced significantly, delivering accurate predictions of human attention.
Figure 1: Saliency and GradCAM identify important regions. Upper row: saliency maps predicted by SOTA model TranSalNet [3]. Bottom row: GradCAM extracted from our IQA model (Baseline-EfficientB0).
@@
  • We propose a methodology to extract saliency maps from trained IQA models, which reveals that learning-based IQA methods incorporate saliency in their predictions and can even outperform saliency prediction baselines, such as center-prior.
@@ II-A Image Quality Assessment
IQA aims to measure the perceptual quality of an image by assessing image artifacts and distortions. Most IQA techniques can be categorized as Full-Reference (FR) or No-Reference (NR), with the latter more relevant for real-world applications [4], where a reference is not available.
Early approaches relied on natural scene statistics (NSS) [5, 6, 7, 8, 9] with hand-crafted features and regression models. As deep learning evolved, neural network-based approaches became popular, typically consisting of a convolutional or transformer backbone with a regression head. HyperIQA [10] proposed a multiscale feature, while MUSIQ [11] employed multiscale inputs and a vision transformer backbone.
Some methods [12, 13] create paired models for both FR and NR scenarios, such as TOPIQ, which uses features from original, distorted, and difference images at each layer.
@@ II-B Saliency in Image Quality Assessment
Research indicates a connection between saliency and IQA [12, 14, 15, 16, 17]. Efforts to enhance IQA models by incorporating saliency include [15, 18, 19, 20]. Early studies [9] improved NSS-based metrics by reweighting error maps based on saliency. SGDNet [18] introduced a separate head to predict saliency and reweight features.
TranSLA [21] incorporates saliency as a query branch in cross-attention, while HVS-5M [19] uses a pre-trained SP model to reweight features. SCVC [22] aggregates patch scores using Gaussian functions and saliency maps (SM). LPIPS [23] performance can be enhanced by spatially reweighting feature maps [12]. However, some studies [14] report only slight improvement in VQA models from using saliency, and our research suggests models already incorporate saliency knowledge into their predictions.
@@ II-C AI Explainability
Neural networks, despite their high performance, remain a black box. Methods have been developed to provide insight into their internal operations. CAM [24] considers a simple case of a CNN with a head consisting of a single linear layer after global average pooling (GAP). GradCAM [25] extends CAM to models with arbitrary heads, using gradients w.r.t. the model’s predictions. Subsequent works [26, 27, 28] propose heuristics for improving the method.
Some researchers have studied benchmarking explanation maps. In [26], they suggest masking important image areas and monitoring changes in model confidence. Work [29] introduced a metric, area over the perturbation curve, as a measure of explanation-map fidelity.
Several works address explainability in IQA. In [13], authors showed that IQA models only require half of an image to make accurate predictions. They divided each image into 12 square patches and examined predictions of the transformer-based IQA model as they masked various combinations. They found that using important regions preserves model quality, while using trivial regions decreases it. We simplify this approach and show that masking pixels based on saliency maps or GradCAM yields comparable results.
@@ III-A Extracting Saliency from IQA models
TABLE I: Performance comparison on SALICON[30] dataset. GradCAM-extracted saliency maps from IQA models outperform center-prior baseline, but are surpassed by SOTA saliency prediction methods. The best results are bold, the second-best are underlined, and the third best are italics.
@@
We started our experiments by testing the hypothesis that ground-truth saliency (e.g. from eye-tracker) correlates with explanation GradCAM maps of IQA models, in other words, if it is possible to extract an approximation for saliency from a trained IQA model. In our experiments we used TOPIQ [12], DBCNN [31] and CLIP-IQA [32]. We also trained our simple baseline model, consisting of a backbone, GAP pooling, and MLP head. We tested two backbones: EfficientNet-B0 [34] and ResNet-50 [35]. To build GradCAMs we used HiResCAM [27] method, known to provably reflect the locations the model used for computation. Additionally, to improve the quality of maps and remove noise, we applied smoothing through augmentations and SVD decompositions of feature maps as recommended in [36]. Every image was passed six times through the model with small rotations. Then feature map from the last channel was channel-wise decomposed via SVD-decomposition and the first main component was taken.
We used straightforward saliency baselines – Center Prior, a Gaussian distribution that approximates the saliency averaged over the dataset – and two State-of-the-Art saliency prediction models [33, 3]. We evaluate the model quality based on common saliency metrics – NSS, SIM, CC, and KLD calculated on the validation split of SALICON dataset [30]. Before calculating metrics, we employed map transforms [37] as a post-processing transformation.
@@ III-B Saliency Masking
Inspired by [13], we investigate the significance of individual image regions on image quality by masking them. The primary aim was to determine which interpretation maps best represent important areas of the image: saliency or GradCAM. For each image, the corresponding explanation map was used, with high values indicating important areas. Subsequently, all pixels in the image whose values did not exceed the threshold were masked.
Masking of different image portions was performed using two strategies: Most Relevant First (MoRF) and Least Relevant First (LeRF), as proposed in [29]. In MoRF, pixels with values above the threshold were masked, preserving trivial regions. Conversely, in LeRF, pixels with values below the threshold were masked, preserving important regions. Perturbations, such as filling with black color or ImageNet mean values, were applied. Before thresholding the explanation maps, Gaussian blur with a large kernel (approximately 101 pixels) and a small sigma was applied to ensure the calculation of thresholds corresponding to all quantiles. Fig. 1 shows examples of images whose regions are filled with black. For the experiment, images from the KonIQ-10k dataset were used. The performance of IQA models on masked images was calculated in terms of Spearman’s Rank Order Correlation Coefficient (SROCC).
Fig. 2 presents the results. A comparison was also made to the method from [13], referred to as ”HalfImage”.
Masking important areas (MoRF) decreases image quality more than masking trivial areas (LeRF). The same observation holds for GradCAM. Interestingly, masking with saliency results in a larger performance gap between LeRF and MoRF regions compared to HalfImage, even though the latter employs knowledge of pretrained IQA models. This finding suggests that saliency maps effectively highlight regions that are crucial for IQA models. Furthermore, it is noted that the behavior of models when masked with GradCAM maps and saliency exhibits significant similarity. This similarity can be interpreted as an existing relationship between the attention of IQA models and human visual attention.
@@ III-C Dual-Task Training
After discovering a close connection between IQA and saliency prediction (SP), we decided to train a model on both tasks simultaneously and determine whether we could do so without decreasing performance. We implemented two approaches to achieve this goal.
In the first approach, termed Baseline-Sal-Loss, we added a small decoder to our baseline model. It consisted of a single convolutional layer with a 1×1 kernel, followed by a sigmoid activation function to integrate saliency loss into the model.
The second approach, called Baseline-GradCAM-Loss, utilizes the GradCAM method as a secondary output, which we incorporated into the model’s loss function to enhance saliency prediction. The GradCAM calculation is as follows:
$L^{c}_{\text{GradCAM}}=\operatorname{ReLU}\left(\sum_{k}\alpha^{c}_{k}A^{k}\right),\quad\alpha^{c}_{k}=\frac{1}{Z}\sum_{i}\sum_{j}\frac{\partial y^{c}}{\partial A^{k}_{ij}}$
where $y^{c}$ is model prediction and $A^{k}$ is $k$-th channel of features from last layer. As IQA model has only one output, $c$ equals 1. To ensure differentiability, we detached $\alpha^{c}_{l}$ from the computational graph, calculating gradients exclusively over the channel maps. This modification is integral to the Baseline-GradCAM-Loss model.
We used 80% of the KonIQ-10k dataset for training and the remaining 20% for testing. Additionally, we evaluated our model on the CLIVE, SALICON, and CAT2000 datasets. Since KonIQ-10k does not provide saliency maps, we generated proxy SM using the MSINet model. Each experiment was repeated 10 times to ensure consistency, with the averaged results presented in Table 1.
We compared the dual-task models with the baseline model trained solely on IQA using the saliency extraction technique, Center Prior, and other State-of-the-Art saliency models. Our findings reveal that both dual-task approaches allow for effective training on IQA and SP tasks concurrently. These approaches maintain the performance of IQA models while significantly improving the results for SP.
@@ III-D Non-uniform-Compression Dataset
We hypothesized that the existing IQA datasets lack sufficient complexity (due to almost-uniform compression and degradation), hindering the ability to detect emerging capabilities in IQA models, utilizing saliency. Consequently, we constructed a new dataset of high-quality images compressed with nonuniform codecs called Saliency-Aware Compressed Images Dataset (SACID). Specifically, we employed the custom codec [41] and sourced 50 images from the CLIC-2021 dataset.
@@
To obtain subjective scores, we conducted pairwise comparisons using over 1,400 assessors from the crowdsourcing platform Subjectify.us. Assessors were presented with 3 verification and 25 random pairs of images with different picture qualities and were asked the following question: “You will be shown sequential pairs of images with different picture quality. For each pair, select the image that has the most acceptable quality for viewing, or note that the quality in this pair is almost the same”.
We evaluated all models from the PYIQA toolbox [42], saliency-aware versions of PSNR and SSIM (EW-PSNR and EW-SSIM [9]), our model variants (Baseline, Baseline-Sal-Loss, and Baseline-GradCAM-Loss), and salient deep-learning models (SGDNet[18], etc.). Models were compared in terms of SROCC, PLCC and fraction of concordant pairs (FracCP). To calculate FracCP, groups corresponding to different images were considered, and in each group, the fraction of ordered pairs was counted. The resulting numbers were averaged across all groups.
@@ IV Discussion
Type Method SROCC $\uparrow$ PLCC $\uparrow$ FracCP $\uparrow$
NSS based MS-SSIM [6] 0.807 0.849 0.507
BRISQUE [8] 0.817 0.833 0.572
PSNR 0.822 0.858 0.505
SSIM [5] 0.835 0.880 0.502
VIF [7] 0.848 0.891 0.526
EW-SSIM [9] 0.850 0.867 0.636
EW-PSNR [9] 0.875 0.893 0.678
FR AHIQ [43] 0.789 0.817 0.546
LPIPS [23] 0.818 0.854 0.512
PieAPP [44] 0.833 0.869 0.535
TOPIQ-FR [12] 0.836 0.848 0.629
DISTS [45] 0.871 0.904 0.589
NR MANIQA [46] 0.752 0.773 0.564
TReS [47] 0.765 0.800 0.559
HyperIQA [10] 0.784 0.802 0.562
SGDNet-None [18] 0.820 0.853 0.651
PaQ2PiQ [48] 0.848 0.858 0.619
SGDNet-Output [18] 0.854 0.879 0.687
Baseline-GradCAM-Loss 0.863 0.865 0.624
CLIP-IQA+ [32] 0.869 0.896 0.598
Baseline 0.871 0.881 0.609
TOPIQ-NR [12] 0.876 0.886 0.629
Baseline-Sal-Loss 0.880 0.886 0.630
MUSIQ [11] 0.888 0.899 0.662
DBCNN [31] 0.901 0.917 0.666
Our experiments reveal a clear link between saliency prediction and IQA. Using GradCAM to extract saliency maps from IQA models shows they implicitly capture human attention better than simple baselines like center-prior. Our dual-task training approach demonstrates that explicitly using saliency maintains or slightly improves IQA performance without adding model complexity.
Masking experiments confirm the significance of salient image regions, as masking them significantly reduces IQA scores. However, directly incorporating saliency into neural IQA methods showed limited gains on our non-uniform compression dataset (SACID). This limitation likely stems from existing IQA datasets being dominated by uniform distortions, restricting models’ ability to leverage saliency cues effectively. Therefore, building datasets with diverse, saliency-driven, and non-uniform distortions is important for further research.